the longer leg of a right triangle is 2 cm less than twice the shorter leg. the hypotenuse is 14 cm less than three times the shorter leg
Answer by stanbon(75887)

the longer leg of a right triangle is 2 cm less than twice the shorter leg. the hypotenuse is 14 cm less than three times the shorter leg
---------
shorter leg: x
longer leg: 2x-2
hypot: 3x-14
------------------------
Use Pythagoras to solve for "x":
x^2 + (2x-2)^2 = 14^2
-----
x^2 + 4x^2 -8x +4 = 196
----
5x^2 - 8x - 192 = 0
-----
x = [8 +- sqrt(8^2 - 4*5*-192)]/10
-----
x = [8 +- sqrt(3904)]/10
-----
x = [8 +- 62.48]/10
-----
x = 7.048 or x = -5.448
-----------------------------
Cheers,
Stan H.
-----------------